Mercury freezes at -40oC.  How many Joules of heat energy would you need to heat 2.0 g of Hg(l) from the freezing point to the boiling point (356.6oC)?  The specific heat of mercury is 0.14 J/(g*oC).



1
Expert's answer
2021-04-12T04:58:36-0400

Solution:

Assume that there are no phase transitions (solid → liquid and liquid → gas) and mercury (Hg) remains liquid in the whole temperature range (from -40°C to 356.6°C).


q = m × C × ΔT

q = m × C × (Tf - Ti),

where:

q = amount of heat energy gained or lost by substance (J)

m = mass of sample (g)

C = specific heat capacity (J °C-1 g-1)

Tf = final temperature (°C)

Ti = initial temperature (°C)

 

Thus:

q = (2.0 g) × (0.14 J °C-1 g-1) × (356.6 - (-40))°C = 111.048 J = 111.05 J

q = 111.05 Joules


Answer: 111.05 Joules of heat energy would you need.